OpenStudy (anonymous):

Please help me? Fan/medal/written testimony Given an exponential function for compounding interest, A(x) = P(.91)x, what is the rate of change?

OpenStudy (anonymous):

@amistre64 @sleepyjess @RedNeckOutLaw @TheSmartOne

OpenStudy (amistre64):

what is the general formula for compounding interest?

OpenStudy (amistre64):

we have a specific example, and need to compare it to the general formula

OpenStudy (anonymous):

I don't understand this question at all. could you go more in depth? @amistre64

OpenStudy (amistre64):

compound interest is generated by taking the interest of the pirincple plus added interest. interest is just a percentage (a rate) of the starting balance called the principal, P for 1 period: interest added is just P(r) giving us a new balance P + Pr, or simpy P(1+r) now for the next period we take the interest of the new balance and add it to itself again: P(1+r) + P(1+r)r = P(1+r)(1+r) = P(1+r)^2 as we continue this process we develop a rule in general: for any given period, x, the balance in the account is determined by the formula: P(1+r)^x

OpenStudy (amistre64):

equating these things we have: P(1+r)^x = P(.91)^x or 1+r = .91 is the best i can imagine unless im reading it incorrectly

OpenStudy (amistre64):

it either that or Pr^x = P(.91)^x and theyve worded it rather poorly

OpenStudy (anonymous):

FLVS sucks for wording. their famous for that lol I can say that by an educated guess, it would be -0.09 because its decreasing?

OpenStudy (amistre64):

that does seem to be a fair assessment of the information given yes -.09

OpenStudy (anonymous):

Thank you ^.^

OpenStudy (amistre64):

youre welcome. dbl chk to make sure its not spose to be P(1.91)^x other than that .... you should be good :)

OpenStudy (anonymous):

wait! it is P(1.91)^x

OpenStudy (anonymous):

@amistre64